### v1

# Decentralized Stochastic Nonconvex Optimization under the Relaxed Smoothness

Abstract
This paper studies decentralized optimization problem $f(\mathbf{x})=\frac{1}{m}\sum_{i=1}^{m}f_{i}(\mathbf{x})$, where each local function has the form of $f_{i}(\mathbf{x})={\mathbb{E}}\left[F(\mathbf{x};{\bm{\xi}}_{i})\right]$ which is $(L_{0},L_{1})$-smooth but possibly nonconvex and the random variable ${\bm{\xi}}_{i}$ follows distribution ${\mathcal{D}}_{i}$. We propose a novel algorithm called decentralized normalized stochastic gradient descent (DNSGD), which can achieve the $\epsilon$-stationary point on each local agent. We present a new framework for analyzing decentralized first-order methods in the relaxed smooth setting, based on the Lyapunov function related to the product of the gradient norm and the consensus error. The analysis shows upper bounds on sample complexity of ${\mathcal{O}}(m^{-1}(L_{f}\sigma^{2}\Delta_{f}\epsilon^{-4}+\sigma^{2}\epsilon^{-2}+L_{f}^{-2}L_{1}^{3}\sigma^{2}\Delta_{f}\epsilon^{-1}+L_{f}^{-2}L_{1}^{2}\sigma^{2}))$ per agent and communication complexity of $\tilde{\mathcal{O}}((L_{f}\epsilon^{-2}+L_{1}\epsilon^{-1})\gamma^{-1/2}\Delta_{f})$, where $L_{f}=L_{0}+L_{1}\zeta$, $\sigma^{2}$ is the variance of the stochastic gradient, $\Delta_{f}$ is the initial optimal function value gap, $\gamma$ is the spectral gap of the network, and $\zeta$ is the degree of the gradient dissimilarity. In the special case of $L_{1}=0$, the above results (nearly) match the lower bounds on decentralized nonconvex optimization in the standard smooth setting. We also conduct numerical experiments to show the empirical superiority of our method.
1 Introduction
This paper considers the decentralized stochastic optimization over a connected network of $m$ nodes. Specifically, we study the minimization problem
$\min_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})=\frac{1}{m}\sum_{i=1}^{m}f_{i}({\bf{x}}),$
where $f_{i}:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is the local function on $i$th agent and has the form of
$f_{i}({\bf{x}})={\mathbb{E}}\left[F_{i}({\bf{x}};{\bm{\xi}}_{i})\right].$
We are interested in the setting that each $f_{i}$ is relaxed smooth [54] but possibly nonconvex and the the random variable ${\bm{\xi}}_{i}$ follows distribution ${\mathcal{D}}_{i}$. The optimization theory and algorithms for relaxed smooth problem have received significant attention in recent years [4, 46, 25, 23, 12, 14, 20, 5, 44, 7, 31, 45, 40, 48, 53, 51, 19, 18, 17], since empirical studies suggest the relaxed smoothness can better characterize the training of neural network models than classical smoothness [54, 6, 8]. However, most existing works focus on centralized scenarios, and how to well address the relaxed smooth optimization problem in decentralized scenarios remains unclear.
The main challenge in decentralized optimization is that each agent can only access the information from itself or its neighbors, which leads to the consensus error among local variables [38]. For smooth optimization, the gradient tracking technique successfully uses the gradient Lipschitz continuity to reduce the consensus error and guarantee the convergence of the algorithm [37, 39]. Additionally, the multi-consensus steps with Chebyshev acceleration [28, 1] can further improve the communication complexity, leading to the optimal decentralized methods for both convex and nonconvex cases [43, 49, 16, 24, 35, 29, 33, 50, 3, 52, 32]. For nonsmooth optimization, the recent works focus on finding the (generalized) stationary points of the Lipschitz continuous objective [41, 27, 21, 22], establishing efficient decentralized stochastic algorithms based on the boundedness of local (sub)gradients.
In this paper, we focus on decentralized optimization under the relaxed smoothness, which is more challenging than above classical settings since neither the local functions nor their gradients enjoy the Lipschitz continuity. We propose a novel algorithm called decentralized normalized stochastic gradient descent (DNSGD) to address above issues. Our method does not include the gradient clipping, avoiding the potential inconsistency in the update modes of local variable arising from consensus errors. We further construct a new Lyapunov function related to the product of the gradient norm and the consensus error, which successfully characterizes the convergence of our decentralized method in the relaxed smooth setting that allows local gradients to be neither bounded nor Lipschitz continuous. Moreover, we conduct empirical studies to show the empirical advantage of our proposed method in practice.
2 Preliminaries
We use the bold lowercase letter with the subscript to represent the local variable, e.g., vector ${\bf{x}}_{i}\in\mathbb{R}^{d}$ is a local variable on the $i$th agent. We use the bold uppercase letter to present the aggregated matrix consisting of local variables from all agents, e.g., $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{m}]^{\top}\in\mathbb{R}^{m\times d}$, where ${\bf{x}}_{i}\in\mathbb{R}^{d}$. We also denote the mean vector by the bold lowercase letter with bar, e.g., $\bar{{\bf{x}}}=\frac{1}{m}{\bf{1}}^{\top}{\bf X}\in{\mathbb{R}}^{1\times d}$, where ${\bf{1}}=[1,\dots,1]\in{\mathbb{R}}^{m}$. In addition, the notation $\|\cdot\|$ represents the Euclidean norm (or the Frobenius) norm for given vector (or matrix). In addition, we use $\bf I$} to present the identity matrix. For the ease of presentation, we allow the input of the function be organized as both column vector or row vector.
We impose following common assumptions for nonconvex optimization under the relaxed smoothness.
Assumption 1 (lower boundedness).
We suppose the objective $f$ is lower bounded, i.e., there exists some $f^{*}\in{\mathbb{R}}$ such that
$f^{*}=\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})>-\infty.$
Assumption 2 (relaxed smoothness).
We suppose each local function $f_{i}$ is $(L_{0},L_{1})$-smooth i.e., there exists some $L_{0},L_{1}\geq 0$ such that each local function $f_{i}$ satisfies
$\left\|\nabla f_{i}({\bf{x}})-\nabla f_{i}({\bf{y}})\right\|\leq\left(L_{0}+L_{1}\left\|\nabla f_{i}({\bf{x}})\right\|\right)\left\|{\bf{x}}-{\bf{y}}\right\|$ (1)
for all ${\bf{x}},{\bf{y}}\in{\mathbb{R}}^{d}$ with $\left\|{\bf{x}}-{\bf{y}}\right\|\leq 1/L_{1}$, where we define $1/L_{1}=+\infty$ if $L_{1}=0$.
Remark 1.
The notion of relaxed smoothness was first introduced by Zhang et al. [54], which requires the second-order differentiability. Here, we consider the weaker condition (Assumption 2) characterized by the gradient, which is provided by Zhang et al. [53].
We also introduce the assumption on gradient dissimilarity to describes the heterogeneity of the local functions, which is widely used in distributed optimization [47, 13, 2, 26].
Assumption 3 (bounded gradient dissimilarity).
We suppose there exists some $\zeta\geq 0$ such that the local gradient satisfies
$\left\|\nabla f_{i}({\bf{x}})-\nabla f({\bf{x}})\right\|\leq\zeta$
for all $i\in[m]$ and ${\bf{x}}\in{\mathbb{R}}^{d}$.
The following proposition shows that the objective $f$ is relaxed smooth if the local functions are relaxed smooth with bounded dissimilarity.
Proposition 1.
Under Assumptions 2 and 3, the objective function $f({\bf{x}})=\frac{1}{m}\sum_{i=1}^{m}f_{i}({\bf{x}})$ is $(L_{f},L_{1})$-smooth with $L_{f}=L_{0}+L_{1}\zeta$, i.e., we have
$\left\|\nabla f({\bf{x}})-\nabla f({\bf{y}})\right\|\leq\left(L_{f}+L_{1}\left\|\nabla f({\bf{x}})\right\|\right)\left\|{\bf{x}}-{\bf{y}}\right\|$ (2)
for all ${\bf{x}},{\bf{y}}\in{\mathbb{R}}^{d}$ with $\left\|{\bf{x}}-{\bf{y}}\right\|\leq 1/L_{1}$.
We are interested in decentralize stochastic first-order methods, which can access the stochastic first-order oracle $\nabla F({\bf{x}};{\bm{\xi}}_{i})$ that satisfies the following assumption.
Assumption 4.
For given $i\in[m]$ and ${\bf{x}}\in{\mathbb{R}}^{d}$, we suppose the $i$th agent can draw ${\bm{\xi}}_{i}\sim{\mathcal{D}}_{i}$ and access the stochastic first-order oracle $\nabla F_{i}({\bf{x}};{\bm{\xi}}_{i})$ which satisfies
${\mathbb{E}}[\nabla F_{i}({\bf{x}};{\bm{\xi}}_{i})]=\nabla f_{i}({\bf{x}})\qquad\text{and}\qquad{\mathbb{E}}\left[\left\|\nabla F_{i}({\bf{x}};{\bm{\xi}}_{i})-\nabla f_{i}({\bf{x}})\right\|^{2}\right]\leq\sigma^{2}$
for some $\sigma>0$.
In decentralized setting, each agent can only exchange information with its neighbors in per communication round. One communication step in decentralized optimization is typically presented by multiplication on the mixing matrix ${\bf W}\in{\mathbb{R}}^{m\times m}$, which satisfies the following standard assumption [42, 16, 50].
Assumption 5.
Let ${\bf W}\in\mathbb{R}^{m\times m}$ be the mixing matrix associated with the connected and undirected network. We suppose (a) the matrix $\bf W$} is symmetric with $w_{i,j}\geq 0$ for all $i,j$, and $w_{i,j}\neq 0$ if and only if the $i$th agent and the $j$th agent are connected or $i=j$; (b) It holds $\mathbf{0}\preceq{\bf W}\preceq{\bf I}$, ${\bf W}^{\top}\mathbf{1}={\bf W}\mathbf{1}=\mathbf{1}$, and $\operatorname{null}({\bf I}-{\bf W})=\operatorname{span}(\mathbf{1})$.
We use $\lambda_{2}({\bf W})$ to present the second largest eigenvalue of the mixing matrix $\bf W$}. Under Assumption 5, we have $\lambda_{2}({\bf W})\in[0,1)$, then we define the spectral gap of $\bf W$} as
$\gamma=1-\lambda_{2}({\bf W}).$
Additionally, we present the multi-consensus steps with Chebyshev acceleration in Algorithm 1 [28, 1], which encourage all local variables to converge to the average as follows.
Proposition 2 (Ye et al. [50, Proposition 1]).
Under Assumption 5, the output of Algorithm 1 holds
$\frac{1}{m}\mathbf{1}^{\top}{\bf Y}^{(K)}=\bar{{\bf{y}}}^{(0)}\quad\text{and}\quad\big{\|}{\bf Y}^{(K)}-\mathbf{1}\bar{{\bf{y}}}^{(0)}\big{\|}\leq c_{1}\big{(}1-c_{2}\sqrt{1-\lambda_{2}({\bf W})}\,\big{)}^{K}\big{\|}{\bf Y}^{(0)}-\mathbf{1}\bar{{\bf{y}}}^{(0)}\big{\|},$
where $\bar{{\bf{y}}}^{(0)}=\frac{1}{m}\mathbf{1}^{\top}{\bf Y}^{(0)}$, $c_{1}=\sqrt{14}$, and $c_{2}=1-1/\sqrt{2}$.
Algorithm 1 ${\rm AccGossip}({\bf Y}^{(0)},{\bf W},K)$
[…]
3:for $k=0,1,\dots,K$ do
4:  ${\bf Y}^{(k+1)}=(1+\eta_{y}){\bf W}{\bf Y}^{(k)}-\eta_{y}{\bf Y}^{(k-1)}$
5:end for
6:Output: ${\bf Y}^{(K)}$
3 The Algorithm and Main Results
Algorithm 2 Decentralized Stochastic Normalized Gradient Descent (DSNGD)
1:Input: ${\bar{\bf{x}}}^{0}\in{\mathbb{R}}^{1\times d}$, ${\bf W}\in{\mathbb{R}}^{m\times m}$, $\eta>0$, $\hat{K},K,T,b\in{\mathbb{N}}$
[…]
5:for $t=0,1,\dots,{T-1}$ do
6:   ${\bf U}^{t}=\begin{bmatrix}\dfrac{{\bf{v}}^{t}_{1}}{\left\|{\bf{v}}^{t}_{1}\right\|}&\dots&\dfrac{{\bf{v}}^{t}_{m}}{\left\|{\bf{v}}^{t}_{m}\right\|}\end{bmatrix}^{\top}$
7:  ${\bf X}^{t+1}={\rm AccGossip}\left({\bf X}^{t}-\eta{\bf U}^{t},{\bf W},K\right)$
8:  independently sample ${\bm{\xi}}_{i,k}^{t+1}\sim{\mathcal{D}}_{i}$ for all $k\in[b]$ at each agent $i$
9:   ${\bf G}^{t+1}=\begin{bmatrix}\displaystyle{\frac{1}{b}\sum_{k=1}^{b}\nabla F_{1}({\bf{x}}_{1}^{t+1};{\bm{\xi}}_{1,k}^{t+1})}&\dots&\displaystyle{\frac{1}{b}\sum_{k=1}^{b}\nabla F_{m}({\bf{x}}_{m}^{t+1};{\bm{\xi}}_{m,k}^{t+1})}\end{bmatrix}^{\top}$
10:   ${\bf V}^{t+1}={\rm AccGossip}\left({\bf V}^{t}+{\bf G}^{t+1}-{\bf G}^{t},{\bf W},K\right)$
11:end for
[…]
13:Output: $\hat{\bf{x}}_{i}$ for each agent $i$
We propose decentralized normalized stochastic gradient descent (DNSGD) in Algorithm 2, which combines the techniques of normalized gradient descent [34, 4, 25, 15], gradient tracking [37, 39], and multi-consensus with Chebyshev acceleration [42, 16, 50]. Here the matrices ${\bf X}^{t}$ and ${\bf V}^{t}$ consists of the local variables and local gradient estimator, i.e., ${\bf X}^{t}=[{\bf{x}}_{1}^{t};\dots;{\bf{x}}_{m}^{t}]^{\top}\in\mathbb{R}^{m\times d}$ and ${\bf V}^{t}=[{\bf{v}}_{1}^{t};\dots;{\bf{v}}_{m}^{t}]^{\top}\in\mathbb{R}^{m\times d}$.
It is worth noting that the relaxed smoothness condition (Assumption 2) does not guarantee that the local gradient is Lipschitz continuous or bounded. However, the existing analysis of decentralized methods mainly depends on the Lipschitz continuity or the boundedness of the local gradient [43, 49, 16, 24, 35, 29, 33, 50, 3, 52, 32, 22, 21, 27, 41, 25], which is not applicable to our setting.
To address above issues, we design the Lyapunov function
$\Phi^{t}=f({\bar{\bf{x}}}^{t})+\frac{3\eta}{\sqrt{m}}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+\frac{2\eta}{\sqrt{m}}\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|,$
where $M_{0}=\sqrt{2(L_{0}^{2}+L_{1}^{2}\zeta^{2})}$ and $M_{1}=\sqrt{2}L_{1}$. In particular, our $\Phi_{t}$ includes the term
$\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|$
to align with the measurement unit on the right-hand side of inequality (2) for the relaxed smoothness of the objective. This design can well address the consensus errors $\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|$ and $\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|$ even if the local gradient is non-Lipschitz and unbounded.
In contrast, the existing analysis essentially only considers the linear combinations of the function value, gradient norm, and the consensus errors [43, 49, 16, 24, 35, 29, 33, 50, 3, 52, 32, 22, 21, 27, 41, 25], so that it does not work on the relaxed smoothness.
By using the Lyapunov function $\Phi^{t}$, we can show the main theoretical results of proposed DSNGD (Algorithm 2) as follows.
Theorem 1.
Under Assumptions 15, running DSNGD (Algorithm 2) with
$\eta=\min\left\{\frac{\epsilon}{4L_{f}+1},\frac{1}{2L_{1}}\right\},\quad b\geq\max\left\{\frac{256(4L_{f}+1)^{2}\sigma^{2}}{mL_{f}^{2}\epsilon^{2}},\frac{1024L_{1}^{2}\sigma^{2}}{mL_{f}^{2}}\right\}$
$T\geq\max\left\{\frac{8(4L_{f}+1)\Delta_{\Phi}}{\epsilon^{2}},\frac{16L_{1}\Delta_{\Phi}}{\epsilon}\right\},\qquad\text{and}\qquad K={\mathcal{O}}\left(\frac{\log m}{\sqrt{\gamma}}\right)$
guarantees ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}})\right\|\right]\leq\epsilon/2$, where $\hat{\bf{x}}$ is uniformly sampled from $\{{\bar{\bf{x}}}^{0},\dots,{\bar{\bf{x}}}^{T-1}\}$,
$L_{f}=L_{0}+L_{1}\zeta,\qquad\text{and}\qquad\Delta_{\Phi}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})+\frac{2\eta}{\sqrt{m}}{\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right].$
Note that the expected $(\epsilon/2)$-stationary point $\hat{\bf{x}}$ in Theorem 1 is based on the mean vectors $\{{\bar{\bf{x}}}^{i}\}_{i=1}^{m}$, which are unavailable in decentralized setting. In addition, the expression of $\Delta_{\Phi}$ indicates the initial consensus error $\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|$ may heavily affect the iteration number $T$ if it is significantly larger than $\Delta_{f}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})$. Fortunately, we can address above issues by appropriate setting of $\hat{K}$ in line 4 to reduce $\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|$ and achieve $\Delta_{\Phi}=\Theta(\Delta_{f})$. Finally, we achieve the refined upper complexity bounds.
Corollary 1.
Following the setting of Theorem 1, we further let $\hat{K}=\tilde{\mathcal{O}}(1/\sqrt{\gamma})$, then it holds that $\Delta_{\Phi}=\Theta(\Delta_{f})$ and the output of DSNGD (Algorithm 2) guarantees ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}}_{i})\right\|\right]\leq\epsilon$ for each agent $i\in[m]$, where $\hat{\bf{x}}_{i}$ is uniformly sampled from $\{{\bf{x}}_{i}^{0},\dots,{\bf{x}}_{i}^{T-1}\}$ generated by the $i$th agent. Additionally, each agent requires the sample complexity of
$Tb={\mathcal{O}}\left(\frac{1}{m}\left(\frac{L_{f}\sigma^{2}\Delta}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}+\frac{L_{1}^{3}\sigma^{2}\Delta}{L_{f}^{2}\epsilon}+\frac{L_{1}^{2}\sigma^{2}}{L_{f}^{2}}\right)\right)$ (3)
and the algorithm requires the communication complexity of
$TK+\hat{K}=\tilde{\mathcal{O}}\left(\frac{L_{f}\Delta_{f}}{\sqrt{\gamma}\epsilon^{2}}+\frac{L_{1}\Delta_{f}}{\sqrt{\gamma}\epsilon}\right),$
[…]
The term $1/m$ in the sample complexity bound (3) implies our method achieves the linear speedup. Although expression (3) looks somewhat complicated, it is easy to understand since we can verify our bounds match or surpass the best-known results in several special cases:
  • In the case of $L_{1}=0$, the local functions reduce to the setting of standard $L_{0}$-smooth and it leads to $L_{f}=L_{0}$, since the objective $f$ is naturally $L_{0}$-smooth without Assumption 3. Then Corollary 1 indicates the sample complexity and the communication complexity of
    $Tb={\mathcal{O}}\left(\frac{1}{m}\left(\frac{L_{0}\sigma^{2}\Delta_{f}}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}\right)\right)\qquad\text{and}\qquad TK+\hat{K}=\tilde{\mathcal{O}}\left(\frac{L_{0}\Delta_{f}}{\gamma^{1/2}\epsilon^{2}}\right),$
    respectively. This matches the (near-)optimal results for decentralized stochastic nonconvex optimization under the standard smoothness [32, Corollary 1].
  • In the case of $m=1$, the problem reduces to the relaxed smooth optimization on single machine. This leads to $\zeta=0$ and $L_{f}=L_{0}$ since there is no heterogeneity. Then Theorem 1 indicates the sample complexity of
    $Tb={\mathcal{O}}\left(\frac{L_{0}\sigma^{2}\Delta}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}+\frac{L_{1}^{3}\sigma^{2}\Delta}{L_{0}^{2}\epsilon}+\frac{L_{1}^{2}\sigma^{2}}{L_{0}^{2}}\right),$ (4)
    which is tighter than the complexity of gradient clipping method in the non-distributed case [53]. Recall that Theorem 3.2 of Zhang et al. [53] claims the sample complexity of ${\mathcal{O}}(L_{0}\sigma^{2}\Delta\epsilon^{-4}+L_{0}^{-3}L_{1}^{4}\sigma^{2}\Delta)$. It is worth noting that term $L_{0}^{-3}L_{1}^{4}\sigma^{2}\Delta$ dominates their bound only when $\epsilon=\Omega(L_{0}/L_{1})$, and it is larger than the term $L_{0}^{-2}L_{1}^{3}\sigma^{2}\Delta\epsilon^{-1}$ in our (4) in this case. In addition, Zhang et al. [53] assume the algorithm can access the exact gradient at the initial point, while their complexity bound does not include the cost of achieving the the exact gradient at the initial point. In contrast, our analysis considers the access of the initial gradient estimator by stochastic first-order oracle, leading to the terms $\sigma^{2}\epsilon^{-2}$ and $L_{1}^{2}L_{0}^{-2}\sigma^{2}$ in our upper bound (4). Recently, Liu & Zhou [30] study the more general stochastic gradient estimator with heavy-tail noise, while applying their result to our setting leads to the term $\sigma^{3}\epsilon^{-3}$ in the upper bound on the sample complexity [30, Theorem 3.2 with $p=2$].
  • Following the case of $m=1$, we further consider the deterministic case, i.e., replace the mini-batch stochastic gradient in Algorithm 2 with the exact gradient and omit the communication steps. Then Theorem 1 indicates the iteration (gradient) complexity of
    $T={\mathcal{O}}\left(\frac{L_{0}\Delta_{f}}{\epsilon^{2}}+\frac{L_{1}\Delta_{f}}{\epsilon}\right),$
    matching the best-known result for solving relaxed smooth problem by exact gradient method on single machine [46, Theorem 3.1].
4 The Complexity Analysis
In this section, we briefly sketch the derivations of our main results. The details for all proofs are presented in appendices. We first present the descent lemma with respect to the mean vectors ${\bar{\bf{x}}}^{t+1}=\frac{1}{m}\sum_{i=1}^{m}{\bf{x}}_{i}^{t+1}$ and ${\bar{\bf{x}}}^{t}=\frac{1}{m}\sum_{i=1}^{m}{\bf{x}}_{i}^{t}$ as follows.
Lemma 1.
Under the setting of Theorem 1, we have
${\mathbb{E}}[f({\bar{\bf{x}}}^{t+1})]\leq$ $f({\bar{\bf{x}}}^{t})-\eta\left(1-\frac{\eta L_{1}}{2}\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|+\frac{2\sqrt{2}\eta}{\sqrt{m}}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|$
$+\frac{\eta}{\sqrt{m}}\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|+\frac{2\eta\sigma}{\sqrt{mb}}+\frac{\eta^{2}L_{f}}{2}$
for given ${\bf X}^{t}$ and ${\bf V}^{t}$.
We define $\rho=c_{1}(1-c_{2}\sqrt{1-\lambda_{2}({\bf W})}\,)^{K}$ with $c_{1}=\sqrt{14}$ and $c_{2}=1-1/\sqrt{2}$. Proposition 2 implies we can guarantee $\rho$ be sufficient small by taking appropriate communication rounds number $K$. We then establish the recursions for the consensus errors as follows.
Lemma 2.
For Algorithm 2, we have
$\left\|{\bf X}^{t+1}-{\bf{1}}{\bar{\bf{x}}}^{t+1}\right\|\leq\rho\left(\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+m\eta\right)\qquad\text{and}\qquad\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|\leq\frac{\rho m\eta}{1-\rho}.$
Lemma 3.
For Algorithm 2 with $K\geq\sqrt{1/\gamma}\log(1+m\eta L_{1})$, we have
$\small\begin{split}&{\mathbb{E}}\left[\left\|{\bf V}^{t+1}-{\bf{1}}{\bar{\bf{v}}}^{t+1}\right\|\right]\\ \leq&\rho\left(\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|+\left(M_{2}+M_{3}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+\left(M_{2}+M_{3}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)m\eta+\frac{2\sqrt{m}\sigma}{\sqrt{b}}\right)\!,\!\!\end{split}$
for given ${\bf X}^{t}$ and ${\bf V}^{t}$, where $M_{2}=(\rho+1)L_{f}+\rho L_{f}L_{1}\eta$ and $M_{3}=\rho L_{1}(L_{1}\eta+1)+L_{1}$.
Applying Lemmas 1, 2, and 3, we can characterize the decease of the Lyapunov function as
${\mathbb{E}}[\Phi^{t+1}]\leq\Phi^{t}-\frac{5\eta}{8}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|+\frac{3\eta^{2}L_{f}}{4}.$ (5)
Summing equation (5) over $t=0,\dots,T-1$, we can achieve our main result presented in Theorem 1 that guarantees ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}})\right\|\right]\leq\epsilon/2$, where $\hat{\bf{x}}$ is uniformly sampled from $\{{\bar{\bf{x}}}^{1},\dots,{\bar{\bf{x}}}^{m}\}$.
In Theorem 1, the setting of iterations number $T\geq\max\left\{8(4L_{f}+1)\Delta_{\Phi}\epsilon^{-2},16L_{1}\Delta_{\Phi}\epsilon^{-1}\right\}$ depends on the term $\Delta_{\Phi}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})+(2\eta/\sqrt{m}\,){\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]$, which may be dominated by the expected consensus error ${\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]$. However, the results in the single machine scenario only depend on $\Delta_{f}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})$. To fill this gap, we set the number of communication rounds at to be $\hat{K}=\tilde{\mathcal{O}}(1/\sqrt{\gamma})$ for line 4 of Algorithm 2, which guarantees that the consensus error satisfies $(2\eta/\sqrt{m}\,){\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]\leq{\mathcal{O}}(\Delta_{f})$ and achieves $\Delta_{\Phi}\leq{\mathcal{O}}(\Delta_{f})$. Therefore, the upper bound of $T$ to guarantee ${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}})\right\|\right]\leq\epsilon/2$ can be improved to ${\mathcal{O}}(L_{f}\Delta_{f}\epsilon^{-2}+L_{1}\Delta_{f}\epsilon^{-1})$.
For achieving an $\epsilon$-stationary point on each node, we consider the gradient at $\hat{\bf{x}}_{i}$ as follows
${\mathbb{E}}[\left\|\nabla f(\hat{\bf{x}}_{i})\right\|]={\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bf{x}}_{i}^{t})\right\|\right]\leq{\mathbb{E}}\bigg{[}\underbrace{\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bf{x}}_{i}^{t})-\nabla f({\bar{\bf{x}}}^{t})\right\|}_{A_{1}}+\underbrace{\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|}_{A_{2}}\bigg{]}.$
For term $A_{1}$, Assumption 2 implies
${\mathbb{E}}[A_{1}]\leq{\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\big{(}L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\big{)}\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|\right]\leq{\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\frac{\rho m\eta\big{(}L_{f}+L_{1}TB\big{)}}{1-\rho}\right],$
where the last step is based on Lemma 8. Recall that Theorem 1 shows ${\mathbb{E}}[A_{2}]={\mathbb{E}}[\left\|\nabla f(\hat{\bf{x}})\right\|]\leq\epsilon/2$, then we can set appropriate $K=\tilde{\mathcal{O}}(\sqrt{1/\gamma}\,)$ to make $\rho$ be sufficient small to guarantee ${\mathbb{E}}[A_{1}]\leq\epsilon/2$. Therefore, we achieve the result of Corollary 1.
5 Discussion and Related Work
The design of our DSNGD (Algorithm 2) is based on normalized (stochastic) gradient descent, which is extensively studied in centralized scenario [34, 4, 9, 15, 11]. Recently, Li et al. [25] studied normalized stochastic gradient descent in decentralized setting, while their analysis is only limited to the setting of standard smoothness. In a very recent work, Jiang et al. [17, Algorithm 2 and Theorem 4] claims a first-order method for decentralized optimization under the relaxed smoothness, while each of its iteration requires every node to perform the step of gradient clipping which depends on accessing the mean vector ${\bar{\bf{x}}}_{i}^{t}$ and the corresponding exact local gradient $\nabla f_{i}({\bar{\bf{x}}}^{t})$. Recall that the decentralized method typically only allows each node directly communicates with its neighbors (Assumption 5) and the stochastic first-order oracle can only access the unbiased gradient estimator with bounded variance (Assumption 4). Therefore, we argue that their algorithm is not truly decentralized or stochastic. In contrast, our method only requires each node to iterate with its local stochastic gradient (line 9 of Algorithm 2), which is fully decentralized and stochastic.
The convergence analysis for our DSNGD (Algorithm 2) depends on the assumption of bounded gradient dissimilarity (Assumption 3), which is commonly used in distributed optimization [47, 13, 2, 26]. Specifically, we use this assumption to establish the inequality $\left\|\nabla f({\bar{\bf{x}}}^{t+1})\right\|\leq(L_{0}+L_{1}\zeta)\eta+\left(1+L_{1}\eta\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|$, which connects the gradients at points gradients at points ${\bar{\bf{x}}}^{t}$ and ${\bar{\bf{x}}}^{t+1}$ (see Lemma 9 in Appendix C). We also combine Assumption 3 with the $(L_{0},L_{1})$-smoothness of local functions (Assumption 2) to guarantee the global objective $f=\frac{1}{m}\sum_{i=1}^{m}f_{i}$ is $(L_{0}+L_{1}\zeta,L_{1})$-smooth (Appendix B). How to establish the convergence for decentralized relaxed smooth optimization without Assumption 3 is still an open problem. It is worth noting that simply assuming that each local function $f_{i}$ is $(L_{0},L_{1})$-smooth even cannot guarantee that the global function $f=\frac{1}{m}\sum_{i=1}^{m}f_{i}$ is also $(L_{0},L_{1})$-smooth, which is different with the standard smooth case. We illustrate this point with the following example.
Proposition 3.
We consider functions
$f_{1}(x)=\exp(Lx),\quad f_{2}(x)=\exp(-Lx),\quad\text{and}\quad f(x)=\frac{f_{1}(x)+f_{2}(x)}{2}$
for some $L>0$ and let $L_{1}=L/\log 2$, then both functions $f_{1}$ and $f_{2}$ is $(0,L_{1})$-smooth while the function $f$ is not $(0,L_{1})$-smooth.
Recall that the discussion after Corollary 1 shows that the sample complexity of our method matches the lower bound in the special case of $L_{1}=0$ [32]. This implies that the ${\mathcal{O}}(\epsilon^{-4})$ dependence on accuracy we obtained cannot be improved in our setting. If we target to achieve an $\mathcal{O}(\epsilon^{-3})$ dependence in the sample complexity, the assumption of expected relaxed smoothness is necessary [4, 40], e.g., strengthen condition (1) in Assumption 2 to ${\mathbb{E}}[\left\|\nabla F_{i}({\bf{x}};{\bm{\xi}}_{i})-\nabla F_{i}({\bf{y}};{\bm{\xi}}_{i})\right\|^{2}]\leq\left(L_{0}+L_{1}\left\|\nabla f_{i}({\bf{x}})\right\|\right)^{2}\left\|{\bf{x}}-{\bf{y}}\right\|^{2}$, then applying the variance-reduced gradient estimators [4, 40] on local agents is possible to achieve the sample complexity depends on $\mathcal{O}(\epsilon^{-3})$.
6 Experiments
In this section, we empirically compare our decentralized normalized stochastic gradient descent (DNSGD) with baseline methods, including decentralized stochastic gradient descent (D-SGD) [26], decentralized stochastic gradient descent with gradient tracking (D-SGT) [36, 39], and decentralized normalized averaged stochastic approximation (D-NASA) [25]. The experiments are performed on a virtualized cluster with 8 vGPUs (NVIDIA RTX 5090, 25.6 GB memory), 8 vCPUs (Intel Xeon Gold 6530), and 96 GB memory.
We present the numerical results on the application of image classification for the MNIST dataset. Specifically we train a three-layer neural network, i.e., the first layer consists of a convolutional layer with one input channel and $16$ output channels, followed by batch normalization, a ReLU activation function, and max pooling; the second layer mirrors the first but has $16$ input channels and $32$ output channels; the last layer is a fully connected layer that predicts $10$ classes.
We conduct the experiments on a distributed system with $10$ agents and its topology is based on the Erdös-Rényi graph [10] with the connectivity probability of $p=0.6$. For all algorithms, we set batch size to $b=256$. The step size of D-SGD, D-SGT and DNSGD are chosen from $\{0.0001,0.0005,0.001,0.005,0.01,0.05,0.1,0.5,1\}$. For our DNSGD, we set the iterations number of Chebyshev acceleration (Algorithm 1) to $\hat{K}=K=2$. We show the results for the gradient norm and the test accuracy in Figures 1 and 2, demonstrating that the proposed DNSGD outperforms all baseline methods.
Refer to caption Refer to caption
(a) sample complexity vs. gradient norm (b) communication rounds vs. gradient norm
Figure 1: We present the sample complexity and the communication rounds against the average of gradient norms $\frac{1}{m}\sum_{i=1}^{m}\nabla f_{i}({\bf{x}}_{i}^{t})$ for MNIST dataset.
Refer to caption Refer to caption
(a) sample complexity vs. test accuracy (b) communication rounds vs. test accuracy
Figure 2: We present the sample complexity and the communication rounds against the test accuracy for MNIST dataset.
7 Conclusion
This paper proposes decentralized normalized stochastic gradient descent for nonconvex optimization under the relaxed smoothness. We provide theoretical analysis to show our method can achieve the $\epsilon$-stationary point on each agent. The empirical studies show that our method performs better than baselines. In future work, we would like to establish decentralized methods for convex optimization under the relaxed smoothness [5, 12, 14, 20, 31, 46]. It is also interesting to extend our results to the more general assumption of symmetric relaxed smoothness [4] and the stochastic gradient estimator with heavily tail-noise [30].
Appendix A Some Technical Lemmas
We first provide some technical lemmas for later proofs.
Lemma 4 (Ye et al. [50, Lemma 12]).
For given matrix ${\bf G}=[{\bf{g}}_{1};\dots;{\bf{g}}_{m}]\in{\mathbb{R}}^{m\times d}$, we have $\left\|{\bf G}-{\bf{1}}{\bar{\bf{g}}}\right\|^{2}\leq\left\|{\bf G}\right\|^{2}$, where ${\bar{\bf{g}}}=\frac{1}{m}\sum_{i=1}^{m}{\bf{g}}_{i}$.
Lemma 5 (Li et al. [25, Lemma B.5]).
For given vectors ${\bf{v}}_{1},\dots,{\bf{v}}_{m}\in{\mathbb{R}}^{1\times d}$ and ${\bar{\bf{v}}}=\frac{1}{m}\sum_{i=1}^{m}{\bf{v}}_{i}$, we have
$\left\|\frac{1}{m}\sum_{i=1}^{m}\frac{{\bf{v}}_{i}}{\left\|{\bf{v}}_{i}\right\|}-\frac{{\bar{\bf{v}}}}{\left\|{\bar{\bf{v}}}\right\|}\right\|\leq\frac{1}{m}\sum_{i=1}^{m}\frac{\left\|{\bar{\bf{v}}}-{\bf{v}}_{i}\right\|}{\left\|{\bar{\bf{v}}}\right\|}.$ (6)
Proof.
This result is presented in the analysis Li et al. [25, Lemma B.5]. Here, we provide the detailed proof for completeness. We have
$\left\|\frac{1}{m}\sum_{i=1}^{m}\frac{{\bf{v}}_{i}}{\left\|{\bf{v}}_{i}\right\|}-\frac{{\bar{\bf{v}}}}{\left\|{\bar{\bf{v}}}\right\|}\right\|$
$=$ $\left\|\frac{1}{m}\sum_{i=1}^{m}\frac{(\left\|{\bar{\bf{v}}}\right\|-\left\|{\bf{v}}_{i}\right\|){\bf{v}}_{i}}{\left\|{\bf{v}}_{i}\right\|\left\|{\bar{\bf{v}}}\right\|}\right\|$
$\leq$ $\frac{1}{m}\sum_{i=1}^{m}\left\|\frac{(\left\|{\bar{\bf{v}}}\right\|-\left\|{\bf{v}}_{i}\right\|){\bf{v}}_{i}}{\left\|{\bf{v}}_{i}\right\|\left\|{\bar{\bf{v}}}\right\|}\right\|$
$\leq$ $\frac{1}{m}\sum_{i=1}^{m}\frac{\left|\left\|{\bar{\bf{v}}}\right\|-\left\|{\bar{\bf{v}}}_{i}\right\|\right|\,\left\|{\bf{v}}_{i}\right\|}{\left\|{\bf{v}}_{i}\right\|\left\|{\bar{\bf{v}}}\right\|}$
$=$ $\frac{1}{m}\sum_{i=1}^{m}\frac{\left|\left\|{\bar{\bf{v}}}\right\|-\left\|{\bf{v}}_{i}\right\|\right|}{\left\|{\bar{\bf{v}}}\right\|}$
$\leq$ $\frac{1}{m}\sum_{i=1}^{m}\frac{\left\|{\bar{\bf{v}}}-{\bf{v}}_{i}\right\|}{\left\|{\bar{\bf{v}}}\right\|},$
where the first inequality is based on vector AM-QM inequality; the second inequality is based on Cauchy–Schwarz inequality; the last step is based on the triangle inequality. ∎
Lemma 6.
The update of Algorithm 2 implies ${\bar{\bf{v}}}^{t}={\bar{\bf{g}}}^{t}$.
Proof.
This result can be obtained by following Luo et al. [33, Lemma 2]. Here, we provide the detailed proof for completeness. The update in line 4 and Proposition 2 imply ${\bar{\bf{v}}}^{0}={\bar{\bf{g}}}^{0}$. Suppose it holds ${\bar{\bf{v}}}^{t}={\bar{\bf{g}}}^{t}$, then the update in line 10 means
${\bar{\bf{v}}}^{t+1}=\frac{1}{m}\sum_{i=1}^{m}({\bf{v}}_{i}^{t}+{\bf{g}}_{i}^{t+1}-{\bf{g}}_{i}^{t})={\bar{\bf{v}}}^{t}+{\bar{\bf{g}}}^{t+1}-{\bar{\bf{g}}}^{t}={\bar{\bf{g}}}^{t+1}.$
Therefore, we finish the proof by induction. ∎
Lemma 7 (Zhang et al. [53, Appendix A.1]).
Under Assumptions 2 and 3, we have
$f({\bf{y}})\leq f({\bf{x}})+\left\langle\nabla f({\bf{x}}),{\bf{y}}-{\bf{x}}\right\rangle+\frac{L_{f}+L_{1}\left\|\nabla f({\bf{x}})\right\|}{2}\left\|{\bf{y}}-{\bf{x}}\right\|^{2}$ (7)
for all ${\bf{x}},{\bf{y}}\in{\mathbb{R}}^{d}$ with $\left\|{\bf{y}}-{\bf{x}}\right\|\leq 1/L_{1}$.
Proof.
Proposition 1 means the objective $f$ is $(L_{f},L_{1})$-smooth, then applying equation (29) of Zhang et al. [53, Appendix A.1] directly achieves equation (7). ∎
Appendix B The Proof of Proposition 1
Proof.
For all ${\bf{x}},{\bf{y}}\in{\mathbb{R}}^{d}$ with $\left\|{\bf{x}}-{\bf{y}}\right\|\leq 1/L_{1}$, we have
$\left\|\nabla f({\bf{x}})-\nabla f({\bf{y}})\right\|^{2}$
$=$ $\left\|\frac{1}{m}\sum_{i=1}^{m}\big{(}\nabla f_{i}({\bf{x}})-\nabla f_{i}({\bf{y}})\big{)}\right\|^{2}$
$\leq$ $\frac{1}{m}\sum_{i=1}^{m}\left\|\nabla f_{i}({\bf{x}})-\nabla f_{i}({\bf{y}})\right\|^{2}$
$\leq$ $\frac{1}{m}\sum_{i=1}^{m}(L_{0}+L_{1}\left\|\nabla f_{i}({\bf{x}})\right\|)^{2}\left\|{\bf{x}}-{\bf{y}}\right\|^{2}$
$\leq$ $(L_{0}+L_{1}(\left\|\nabla f_{i}({\bf{x}})-\nabla f({\bf{x}})\right\|+\left\|\nabla f({\bf{x}})\right\|))^{2}\left\|{\bf{x}}-{\bf{y}}\right\|^{2}$
$=$ $(L_{0}+L_{1}\zeta+L_{1}\left\|\nabla f({\bf{x}})\right\|)^{2}\left\|{\bf{x}}-{\bf{y}}\right\|^{2},$
where the first inequality is based on the Young’s inequality; the second inequality is based on Assumption 2; the third inequality is based on Assumption 3. Therefore, the function $f$ is $(L_{f},L_{1})$-smooth with $L_{f}=L_{0}+L_{1}\zeta$. ∎
Appendix C The Proof of Lemma 1
We first provide the upper bound for the difference of mean vectors ${\bar{\bf{x}}}^{t+1}$ and ${\bar{\bf{x}}}^{t}$.
Lemma 8.
Under the setting of Theorem 1, we have $\left\|{\bar{\bf{x}}}^{t+1}-{\bar{\bf{x}}}^{t}\right\|^{2}\leq\eta^{2}\leq 1/(4L_{1}^{2})$.
Proof.
The updates in lines 6 and 7 of Algorithm 2 imply
${\bar{\bf{x}}}^{t+1}=$ ${\bar{\bf{x}}}^{t}-\eta{\bar{\bf{u}}}^{t}={\bar{\bf{x}}}^{t}-\eta\cdot\frac{1}{m}\sum_{i=1}^{m}\frac{{\bf{v}}^{t}_{i}}{\left\|{\bf{v}}^{t}_{i}\right\|},$ (8)
where the first step is based on Proposition 2. Therefore, we have
$\begin{split}\left\|{\bar{\bf{x}}}^{t+1}-{\bar{\bf{x}}}^{t}\right\|^{2}=\eta^{2}\left\|\frac{1}{m}\sum_{i=1}^{m}\frac{{\bf{v}}^{t}_{i}}{\left\|{\bf{v}}^{t}_{i}\right\|}\right\|^{2}\leq\frac{\eta^{2}}{m}\sum_{i=1}^{m}\left\|\frac{{\bf{v}}^{t}_{i}}{\left\|{\bf{v}}^{t}_{i}\right\|}\right\|^{2}=\eta^{2}\leq\frac{1}{4L_{1}^{2}},\end{split}$ (9)
where the first inequality is based on the vector AM-QM inequality and the last step is based on the setting of $\eta=\min\left\{\epsilon/(4L_{f}+1),1/(2L_{1})\right\}$. ∎
We then connect the gradient $\nabla f({\bar{\bf{x}}}^{t})$ to $\nabla f({\bar{\bf{x}}}^{t+1})$ as follows.
Lemma 9.
Under the setting of Theorem 1, we have
$\left\|\nabla f({\bar{\bf{x}}}^{t+1})\right\|\leq L_{f}\eta+\left(1+L_{1}\eta\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|.$
Proof.
We have
$\begin{split}&\left\|\nabla f({\bar{\bf{x}}}^{t+1})\right\|\\ \leq&\left\|\nabla f({\bar{\bf{x}}}^{t+1})-\nabla f({\bar{\bf{x}}}^{t})\right\|+\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\\ \leq&\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bar{\bf{x}}}^{t+1}-{\bar{\bf{x}}}^{t}\right\|+\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\\ \leq&\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\eta+\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\\ =&L_{f}\eta+\left(1+L_{1}\eta\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|,\end{split}$ (10)
where the first step is based on triangle inequality; the second step is based on Proposition 1; and third last step is based on Lemma 8. ∎
Now we present the proof of Lemma 1.
Proof.
Applying Lemmas 7, 8 and 9, we have
[…]
where the first inequality is based on Lemma 7 and the fact $\left\|{\bar{\bf{x}}}^{t+1}-{\bar{\bf{x}}}^{t}\right\|\leq 1/(2L_{1})$ from Lemma 8; the forth to last line is based on Cauchy–Schwarz inequality; the third to last line is based on the triangle inequality; the last line is based on the vector AM-QM inequality.
We bound the term $\left\|\nabla f({\bar{\bf{x}}}^{t})-{\bar{\bf{v}}}^{t}\right\|$ in equation (11) as
$\begin{split}&\left\|\nabla f({\bar{\bf{x}}}^{t})-{\bar{\bf{v}}}^{t}\right\|^{2}=\left\|\frac{1}{m}\sum_{i=1}^{m}(\nabla f_{i}({\bar{\bf{x}}}^{t})-{\bf{g}}_{i}^{t}))\right\|^{2}\\ \leq&2\underbrace{\left\|\frac{1}{m}\sum_{i=1}^{m}(\nabla f_{i}({\bar{\bf{x}}}^{t})-\nabla f_{i}({\bf{x}}_{i}^{t}))\right\|^{2}}_{B_{1}}+2\underbrace{\left\|\frac{1}{m}\sum_{i=1}^{m}(\nabla f_{i}({\bf{x}}_{i}^{t})-{\bf{g}}_{i}^{t})\right\|^{2}}_{B_{2}}.\end{split}$ (12)
We bound term $B_{1}$ as
$\begin{split}B_{1}=&\left\|\frac{1}{m}\sum_{i=1}^{m}(\nabla f_{i}({\bar{\bf{x}}}^{t})-\nabla f_{i}({\bf{x}}_{i}^{t}))\right\|^{2}\\ \leq&\frac{1}{m}\sum_{i=1}^{m}\left\|\nabla f_{i}({\bar{\bf{x}}}^{t})-\nabla f_{i}({\bf{x}}_{i}^{t})\right\|^{2}\\ \leq&\frac{1}{m}\sum_{i=1}^{m}2\left(L_{0}^{2}+L_{1}^{2}\left\|\nabla f_{i}({\bar{\bf{x}}}^{t})\right\|^{2}\right)\left\|{\bar{\bf{x}}}^{t}-{\bf{x}}_{i}^{t}\right\|^{2}\\ \leq&\frac{1}{m}\sum_{i=1}^{m}2\left(L_{0}^{2}+L_{1}^{2}\left(\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|+\zeta\right)^{2}\right)\left\|{\bar{\bf{x}}}^{t}-{\bf{x}}_{i}^{t}\right\|^{2}\\ \leq&\frac{1}{m}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)^{2}\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|^{2},\end{split}$ (13)
where we denote $M_{0}=\sqrt{2(L_{0}^{2}+L_{1}^{2}\zeta^{2})}$ and $M_{1}=\sqrt{2}L_{1}$. Here, the first inequality is based on vector AM-QM inequality; the second inequality is based on Young’s inequality and Assumption 2; the third inequality is based on Assumption 3.
For given $\{{\bf{x}}_{i}^{t}\}_{i=1}^{m}$, we bound expectation of term $B_{2}$ as
$\begin{split}{\mathbb{E}}[B_{2}]=&{\mathbb{E}}\left\|\frac{1}{m}\sum_{i=1}^{m}({\bf{g}}_{i}^{t}-\nabla f_{i}({\bf{x}}_{i}^{t}))\right\|^{2}\\ =&\frac{1}{m^{2}}\sum_{i=1}^{m}{\mathbb{E}}\left\|{\bf{g}}_{i}^{t}-\nabla f_{i}({\bf{x}}_{i}^{t})\right\|^{2}\\ \leq&\frac{1}{m^{2}}\sum_{i=1}^{m}\frac{\sigma^{2}}{b}=\frac{\sigma^{2}}{mb},\end{split}$ (14)
where the inequality is based on the setting ${\bf{g}}_{i}^{t}=\frac{1}{b}\sum_{k=1}^{b}\nabla F_{i}({\bf{x}}_{i}^{t+1};{\bm{\xi}}_{i,k}^{t+1})$ in line 9 of Algorithm 2 and Assumption 4.
Combining results (12), (13) and (14), we have
$\begin{split}{\mathbb{E}}\left\|\nabla f({\bar{\bf{x}}}^{t})-{\bar{\bf{v}}}^{t}\right\|\leq&\sqrt{\frac{2}{m}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)^{2}\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|^{2}+\frac{\sigma^{2}}{mb}}\\ \leq&\sqrt{\frac{2}{m}}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+\frac{\sigma}{\sqrt{mb}},\end{split}$ (15)
for given ${\bf X}^{t}$, where the first step is based on the Jensen’s inequality and the second step is based on the elementary inequality $\sqrt{a^{2}+b^{2}}\leq a+b$ for $a,b>0$.
Combining equations (11) and (15), we finish the proof. ∎
Appendix D The Proof of Lemma 2
Proof.
Recall that line 6 of Algorithm 2 set
${\bf U}^{t}=\begin{bmatrix}\dfrac{{\bf{v}}^{t}_{1}}{\left\|{\bf{v}}^{t}_{1}\right\|}&\dots&\dfrac{{\bf{v}}^{t}_{1}}{\left\|{\bf{v}}^{t}_{m}\right\|}\end{bmatrix}^{\top}$,
which implies $\left\|{\bf U}^{t}\right\|^{2}=m$. Based on equation (8), we have
$\begin{split}&\left\|{\bf X}^{t+1}-{\bf{1}}{\bar{\bf{x}}}^{t+1}\right\|\\ =&\left\|{\rm AccGossip}({\bf X}^{t}-\eta{\bf U}^{t},K)-\frac{1}{m}{\bf{1}}{\bf{1}}^{\top}{\rm AccGossip}({\bf X}^{t}-\eta{\bf U}^{t},K)\right\|\\ \leq&\rho\left\|{\bf X}^{t}-\eta{\bf U}^{t}-\frac{1}{m}{\bf{1}}{\bf{1}}^{\top}({\bf X}^{t}-\eta{\bf U}^{t})\right\|\\ =&\rho\left\|{\bf X}^{t}-\eta{\bf U}^{t}-{\bf{1}}({\bar{\bf{x}}}^{t}-\eta{\bar{\bf{u}}}^{t})\right\|\\ \leq&\rho\left(\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+\eta\left\|{\bf U}^{t}-{\bf{1}}{\bar{\bf{u}}}^{t}\right\|\right)\\ \leq&\rho\left(\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+m\eta\right),\end{split}$ (16)
where the first inequality is based on Proposition 2; the second inequality is based on the triangle inequality; the last step applies Lemma 4 which leads to
$\left\|{\bf U}^{t}-{\bf{1}}{\bar{\bf{u}}}^{t}\right\|^{2}\leq\left\|{\bf U}^{t}\right\|^{2}=m.$
Based on equation (16), we have
$\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|\leq$ $\rho^{t}\left(\left\|{\bf X}^{0}-{\bf{1}}{\bar{\bf{x}}}^{0}\right\|-\frac{\rho m\eta}{1-\rho}\right)+\frac{\rho m\eta}{1-\rho}\leq\frac{\rho m\eta}{1-\rho},$
where the first inequality is based on the initialization such that ${\bf{x}}_{i}^{0}=\bar{\bf{x}}^{0}$ for each $i$.
Appendix E The Proof of Lemma 3
Proof.
We denote
$\nabla{\bf{f}}({\bf X}^{t})=\begin{bmatrix}\nabla f_{1}({\bf{x}}_{1}^{t})^{\top}\\ \vdots\\ \nabla f_{m}({\bf{x}}_{m}^{t})^{\top}\\ \end{bmatrix}\in{\mathbb{R}}^{m\times d}\qquad\text{and}\qquad\nabla{\bf{f}}({\bf X}^{t+1})=\begin{bmatrix}\nabla f_{1}({\bf{x}}_{1}^{t+1})^{\top}\\ \vdots\\ \nabla f_{m}({\bf{x}}_{m}^{t+1})^{\top}\\ \end{bmatrix}\in{\mathbb{R}}^{m\times d}.$
The update in line 10 of Algorithm 2 implies
$\small\begin{split}&\left\|{\bf V}^{t+1}-{\bf{1}}{\bar{\bf{v}}}^{t+1}\right\|\\ =&\left\|{\rm AccGossip}({\bf V}^{t}+{\bf G}^{t+1}-{\bf G}^{t}),K)-\frac{1}{m}{\bf{1}}{\bf{1}}^{\top}{\rm AccGossip}({\bf V}^{t}+{\bf G}^{t+1}-{\bf G}^{t}),K)\right\|\\ \leq&\rho\left\|{\bf V}^{t}+{\bf G}^{t+1}-{\bf G}^{t}-\frac{1}{m}{\bf{1}}{\bf{1}}^{\top}({\bf V}^{t}+{\bf G}^{t+1}-{\bf G}^{t})\right\|\\ \leq&\rho\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|+\rho\left\|{\bf G}^{t+1}-{\bf G}^{t}-\frac{1}{m}{\bf{1}}{\bf{1}}^{\top}({\bf G}^{t+1}-{\bf G}^{t})\right\|\\ \leq&\rho\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|+\rho\left\|{\bf G}^{t+1}-{\bf G}^{t}\right\|\\ \leq&\rho\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|+\rho\underbrace{\left\|\nabla{\bf{f}}({\bf X}^{t+1})-\nabla{\bf{f}}({\bf X}^{t})\right\|}_{C_{1}}+\rho\underbrace{\left\|{\bf G}^{t+1}-\nabla{\bf{f}}({\bf X}^{t+1})\right\|}_{C_{2}}+\rho\underbrace{\left\|{\bf G}^{t}-\nabla{\bf{f}}({\bf X}^{t})\right\|}_{C_{3}},\end{split}$ (17)
where the first three step follows the derivation of equation (16); the fourth step is based on Lemma 4; the last step is based on triangle inequality.
We first consider term $C_{1}1$. The setting of $K$ in this lemma implies
$\rho\leq\frac{1}{1+m\eta L_{1}},$
which leads to
$\left\|{\bf{x}}_{t}^{i}-{\bar{\bf{x}}}^{t}\right\|\leq\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|\leq\frac{\rho m\eta}{1-\rho}\leq\frac{1}{L_{1}}.$ (18)
Therefore, we have
$\left\|\nabla f({\bf{x}}^{t+1}_{i})-\nabla f({\bf{x}}^{t}_{i})\right\|$
$\leq$ $\left\|\nabla f({\bf{x}}^{t+1}_{i})-\nabla f({\bar{\bf{x}}}^{t+1})\right\|+\left\|\nabla f({\bar{\bf{x}}}^{t+1})-\nabla f({\bar{\bf{x}}}^{t})\right\|+\left\|\nabla f({\bar{\bf{x}}}^{t}_{i})-\nabla f({\bf{x}}^{t}_{i})\right\|$
$\leq$ $\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t+1})\right\|\right)\left\|{\bf{x}}^{t+1}_{i}-{\bar{\bf{x}}}^{t+1}\right\|+\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bar{\bf{x}}}^{t+1}-{\bar{\bf{x}}}^{t}\right\|$
$+\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf{x}}^{t}_{i}-{\bar{\bf{x}}}^{t}\right\|,$
where the first step is based on the triangle inequality and the last step is based on Proposition 1 and equation (18). Summing over above result with $i=1,\dots,m$, we have
$\small\begin{split}C_{1}=&\sum_{i=1}^{m}\left\|\nabla f({\bf{x}}^{t+1}_{i})-\nabla f({\bf{x}}^{t}_{i})\right\|^{2}\\ =&3\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t+1})\right\|\right)^{2}\sum_{i=1}^{m}\left\|{\bf{x}}^{t+1}_{i}-{\bar{\bf{x}}}^{t+1}\right\|^{2}\\ &+3\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)^{2}\sum_{i=1}^{m}\left\|{\bar{\bf{x}}}^{t+1}-{\bar{\bf{x}}}^{t}\right\|^{2}+3\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)^{2}\sum_{i=1}^{m}\left\|{\bf{x}}^{t}_{i}-{\bar{\bf{x}}}^{t}\right\|^{2}\\ \leq&3\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t+1})\right\|\right)^{2}\left\|{\bf X}^{t+1}-{\bf{1}}{\bar{\bf{x}}}^{t+1}\right\|^{2}\\ &+3m\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)^{2}\eta^{2}+3\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)^{2}\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|^{2},\end{split}$ (19)
where the last step is based on Lemma 8.
Substituting equations (16) and (10) into equation (19), we have
$\begin{split}&\left\|\nabla{\bf{f}}({\bf X}^{t+1})-\nabla{\bf{f}}({\bf X}^{t})\right\|^{2}\\ \leq&3\rho^{2}\left(L_{f}+L_{1}\left(L_{f}\eta+(L_{1}\eta+1)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\right)^{2}\left(\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+m\eta\right)^{2}\\ &+3\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)^{2}\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|^{2}+3m\left(L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)^{2}\eta^{2}\\ \leq&\left(M_{2}+M_{3}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)^{2}\left(\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+m\eta\right)^{2},\end{split}$ (20)
where we take $M_{2}=(\rho+1)L_{f}+\rho L_{f}L_{1}\eta$ and $M_{3}=\rho L_{1}(L_{1}\eta+1)+L_{1}$.
For the term $C_{2}$, we have
$\begin{split}&{\mathbb{E}}\left[\left\|{\bf G}^{t+1}-\nabla{\bf{f}}({\bf X}^{t+1})\right\|^{2}\right]\\ =&{\mathbb{E}}\left[\sum_{i=1}^{m}\left\|{\bf{g}}_{i}^{t+1}-\nabla f_{i}({\bf{x}}_{i}^{t+1})\right\|^{2}\right]\\ =&{\mathbb{E}}\left[\sum_{i=1}^{m}\left\|\frac{1}{b}\sum_{k=1}^{b}\nabla F_{i}({\bf{x}}_{i}^{t+1};{\bm{\xi}}_{i,k}^{t+1})-\nabla f_{i}({\bf{x}}_{i}^{t+1})\right\|^{2}\right]\\ \leq&\frac{m\sigma^{2}}{b},\end{split}$ (21)
for given $\{{\bf{x}}_{i}^{t+1}\}_{i=1}^{m}$, where the inequality is based on Assumption 4. Consequently, we have
[…]
where the first inequality is based on Jensen’s inequality.
Similarly, the term $C_{3}$ holds
$C_{3}={\mathbb{E}}\left[\left\|{\bf G}^{t}-\nabla{\bf{f}}({\bf X}^{t})\right\|\right]\leq\frac{\sqrt{m}\sigma}{\sqrt{b}}.$ (23)
Combining the results of (17), (20), (22), and (23), we have
$\small\begin{split}&{\mathbb{E}}\left[\left\|{\bf V}^{t+1}-{\bf{1}}{\bar{\bf{v}}}^{t+1}\right\|\right]\\ \leq&\rho\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|+\rho\left(M_{2}+M_{3}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left(\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+m\eta\right)+\frac{2\rho\sqrt{m}\sigma}{b}\\ \leq&\rho\left(\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|+\left(M_{2}+M_{3}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+\left(M_{2}+M_{3}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)m\eta+\frac{2\sqrt{m}\sigma}{\sqrt{b}}\right),\end{split}$
for given $\{{\bf{x}}_{i}^{t+1}\}_{i=1}^{m}$ and $\{{\bf{v}}_{i}^{t+1}\}_{i=1}^{m}$, which finishes the proof. ∎
Appendix F The proof of Theorem 1
For given ${\bf X}^{t}$ and ${\bf V}^{t}$, we apply Lemmas 1, 2, 3, and 9 to achieve
${\mathbb{E}}[\Phi^{t+1}]$
$=$ ${\mathbb{E}}\left[f({\bar{\bf{x}}}^{t+1})+\frac{3\eta}{\sqrt{m}}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{t+1})\right\|\right)\left\|{\bf X}^{t+1}-{\bf{1}}{\bar{\bf{x}}}^{t+1}\right\|+\frac{2\eta}{\sqrt{m}}\left\|{\bf V}^{t+1}-{\bf{1}}{\bar{\bf{v}}}^{t+1}\right\|\right]$
$\leq$ $f({\bar{\bf{x}}}^{t})-\eta\left(1-\frac{\eta L_{1}}{2}\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|$
$+\frac{2\sqrt{2}\eta}{\sqrt{m}}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+\frac{\eta}{\sqrt{m}}\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|+\frac{2\eta\sigma}{\sqrt{mb}}+\frac{\eta^{2}L_{f}}{2}$
$+\frac{3\eta\rho}{\sqrt{m}}\left(M_{0}+M_{1}\left(L_{f}\eta+\left(1+L_{1}\eta\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\right)\left(\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+m\eta\right)$
$+\frac{2\eta\rho}{\sqrt{m}}\left(\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|+\left(M_{2}+M_{3}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+\left(M_{2}+M_{3}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)m\eta+\frac{2\sqrt{m}\sigma}{\sqrt{b}}\right)$
$=$ $f({\bar{\bf{x}}}^{t})+\left(\frac{2\sqrt{2}M_{0}\eta}{\sqrt{m}}+\rho\Big{(}\frac{3\eta(M_{0}+M_{1}L_{f}\eta)}{\sqrt{m}}+\frac{2\eta M_{2}}{\sqrt{m}}\Big{)}\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|$
$+\left(\frac{2\sqrt{2}M_{1}\eta}{\sqrt{m}}+\rho\left(\frac{3\eta M_{1}(1+L_{1}\eta)}{\sqrt{m}}+\frac{2\eta M_{3}}{\sqrt{m}}\right)\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+\left(\frac{\eta}{\sqrt{m}}+\frac{2\rho\eta}{\sqrt{m}}\right)\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|$
$-\eta\left(1-\frac{\eta L_{1}}{2}-\rho\Big{(}3\sqrt{m}\eta M_{1}(1+L_{1}\eta)+2\sqrt{m}\eta M_{3}\Big{)}\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|$
$+\frac{2\eta\sigma}{\sqrt{mb}}+\frac{\eta^{2}L_{f}}{2}+\rho\Big{(}3\sqrt{m}\eta^{2}(M_{0}+M_{1}L_{f}\eta)+2\sqrt{m}\eta^{2}M_{2}\Big{)}$
$\leq$ $f({\bar{\bf{x}}}^{t})+\frac{3\eta}{\sqrt{m}}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right)\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|+\frac{2\eta}{\sqrt{m}}\left\|{\bf V}^{t}-{\bf{1}}{\bar{\bf{v}}}^{t}\right\|$
$-\eta\left(\frac{7}{8}-\frac{\eta L_{1}}{2}\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|+\frac{3\eta^{2}L_{f}}{4}$
$=$ $\Phi^{t}-\eta\left(\frac{7}{8}-\frac{\eta L_{1}}{2}\right)\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|+\frac{3\eta^{2}L_{f}}{4}$
$\leq$ $\Phi^{t}-\frac{5\eta}{8}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|+\frac{3\eta^{2}L_{f}}{4},$
where the second inequality requires the condition
$\begin{cases}2\sqrt{2}M_{0}+\rho(3(M_{0}+M_{1}L_{f}\eta)+2M_{2})\leq 3M_{0},\\[2.84544pt] 2\sqrt{2}M_{1}+\rho\left(3M_{1}(1+L_{1}\eta)+2M_{3}\right)\leq 3M_{1},\\[2.84544pt] \eta+2\rho\eta\leq 2\eta,\\[4.26773pt] \rho\Big{(}3\sqrt{m}\eta M_{1}(1+L_{1}\eta)+2\sqrt{m}\eta M_{3}\Big{)}\leq\dfrac{1}{8},\\[7.11317pt] \dfrac{2\sigma}{\sqrt{mb}}+\rho\Big{(}3\sqrt{m}\eta(M_{0}+M_{1}L_{f}\eta)+2\sqrt{m}\eta M_{2}\Big{)}\leq\dfrac{\eta L_{f}}{4}\end{cases}$
that holds by taking $b\geq\left\lceil{256\sigma^{2}}/(m\eta^{2}L_{f}^{2})\right\rceil$ and
$\rho\leq\min\Bigg{\{}$ $\frac{1}{1+m\eta L_{1}},\frac{(3-2\sqrt{2})M_{0}}{3(M_{0}+M_{1}L_{f}\eta)+2M_{2}},\frac{(3-2\sqrt{2})M_{1}}{3M_{1}(1+L_{1}\eta)+2M_{3}},\frac{1}{2},$
$\dfrac{1}{8(3\sqrt{m}\eta M_{1}(1+L_{1}\eta)+2\sqrt{m}\eta M_{3})},\dfrac{\eta L_{f}}{8(3\sqrt{m}\eta(M_{0}+M_{1}L_{f}\eta)+2\sqrt{m}\eta M_{2})}\Bigg{\}}.$
Recall that we have $L_{1}\eta\leq 1/2$, $M_{0}=\Theta(L_{f})$, $M_{1}=\Theta(L_{1})$, $M_{2}\leq\Theta(L_{f})$, and $M_{3}\leq\Theta(L_{1})$, which means we require
$\rho\leq{\mathcal{O}}\left(\frac{1}{\sqrt{m}}\right).$
This can be achieved by taking
$K={\mathcal{O}}\left(\frac{\log m}{\sqrt{\gamma}}\right).$
Following the setting $\eta=\min\left\{\epsilon/(4L_{f}+1),1/(2L_{1})\right\}$, we have
${\mathbb{E}}\left[\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right]\leq\frac{{\mathbb{E}}[8(\Phi^{t}-\Phi^{t+1})]}{5\eta}+\frac{6\eta L_{f}}{5}.$
Taking the average on above inequalities with $t=0,\dots,T-1$, we have
${\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right]\leq{\mathbb{E}}\left[\frac{8(\Phi^{0}-\Phi^{T})}{5\eta T}+\frac{6\eta L_{f}}{5}\right]\leq\frac{8\Delta_{\Phi}}{5\eta T}+\frac{6\eta L_{f}}{5},$
where the last step holds due to
${\mathbb{E}}[\Phi^{0}-\Phi^{T}]$
$=$ ${\mathbb{E}}\Bigg{[}f({\bar{\bf{x}}}^{0})+\frac{3\eta}{\sqrt{m}}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{0})\right\|\right)\left\|{\bf X}^{0}-{\bf{1}}{\bar{\bf{x}}}^{0}\right\|+\frac{2\eta}{\sqrt{m}}\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|$
$-\left(f({\bar{\bf{x}}}^{T})+\frac{3\eta}{\sqrt{m}}\left(M_{0}+M_{1}\left\|\nabla f({\bar{\bf{x}}}^{T})\right\|\right)\left\|{\bf X}^{T}-{\bf{1}}{\bar{\bf{x}}}^{T}\right\|+\frac{2\eta}{\sqrt{m}}\left\|{\bf V}^{T}-{\bf{1}}{\bar{\bf{v}}}^{T}\right\|\right)\Bigg{]}$
$\leq$ $f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})+\frac{2\eta}{\sqrt{m}}{\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]=\Delta_{\Phi}.$
We consider the following case for the choice of $\eta$:
  1. 1.
    In the case of $\epsilon/(4L_{f}+1)\leq 1/(2L_{1})$, we follow the setting
    $\eta=\frac{\epsilon}{4L_{f}+1},\quad b\geq\left\lceil\frac{256(4L_{f}+1)^{2}\sigma^{2}}{mL_{f}^{2}\epsilon^{2}}\right\rceil,\quad\text{and}\quad T\geq\frac{8\Delta_{\Phi}}{\eta\epsilon}=\frac{8(4L_{f}+1)\Delta_{\Phi}}{\epsilon^{2}}$ (24)
    to obtain
    ${\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right]\leq\frac{\epsilon}{5}+\frac{3\epsilon}{10}=\frac{\epsilon}{2}.$
  2. 2.
    In the case of $1/(2L_{1})\leq\epsilon/(4L_{f}+1)$, we follow the setting
    $\eta=\frac{1}{2L_{1}},\qquad b\geq\left\lceil\frac{1024L_{1}^{2}\sigma^{2}}{mL_{f}^{2}}\right\rceil,\qquad\text{and}\qquad T\geq\frac{16L_{1}\Delta_{\Phi}}{\epsilon}.$ (25)
    to obtain
    ${\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right]\leq\frac{\epsilon}{5}+\frac{3\epsilon}{10}=\frac{\epsilon}{2}.$
Hence, we finish the proof.
Appendix G The Proof of Corollary 1
Proof.
We first provide the upper bound for ${\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]$. Let ${\bar{\bf{g}}}^{0}=\frac{1}{m}{\bf{1}}^{\top}{\bf G}^{0}$, then we have
${\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|^{2}\right]\leq{\mathbb{E}}\left[\hat{\rho}^{2}\left\|{\bf G}^{0}-{\bf{1}}{\bar{\bf{g}}}^{0}\right\|^{2}\right]$
$\leq$ ${\mathbb{E}}\left[\hat{\rho}^{2}\left\|{\bf G}^{0}\right\|^{2}\right]=\hat{\rho}^{2}\sum_{i=1}^{m}{\mathbb{E}}\left\|\frac{1}{b}\sum_{k=1}^{b}\nabla F_{i}({\bf{x}}_{i}^{0};{\bm{\xi}}_{i,k}^{0})\right\|^{2}$
$=$ $2\hat{\rho}^{2}\sum_{i=1}^{m}\left({\mathbb{E}}\left[\left\|\frac{1}{b}\sum_{k=1}^{b}\nabla F_{i}({\bf{x}}_{i}^{0};{\bm{\xi}}_{i,k}^{0})-\nabla f_{i}({\bf{x}}_{i}^{0})\right\|^{2}+\left\|\nabla f_{i}({\bf{x}}_{i}^{0})\right\|^{2}\right]\right)$
$=$ $2\hat{\rho}^{2}\sum_{i=1}^{m}\left(\frac{\sigma^{2}}{b}+\left\|\nabla f_{i}({\bf{x}}_{i}^{0})\right\|^{2}\right)$
$=$ $2\hat{\rho}^{2}\left(\frac{m\sigma^{2}}{b}+\sum_{i=1}^{m}\left\|\nabla f_{i}({\bf{x}}_{i}^{0})\right\|^{2}\right).$
Hence, we require
$\hat{\rho}\leq\frac{\sqrt{m}\Delta_{f}}{2\sqrt{2}\eta\sqrt{{m\sigma^{2}}/{b}+\sum_{i=1}^{m}\left\|\nabla f_{i}({\bf{x}}_{i}^{0})\right\|^{2}}}$
to guarantee
${\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]\leq\sqrt{{\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|^{2}\right]}\leq\frac{\sqrt{m}\Delta_{f}}{2\eta},$ (26)
where the first step is based on Jensen’s inequality. Consequently, we have
$\Delta_{\Phi}=f({\bar{\bf{x}}}^{0})-\inf_{{\bf{x}}\in{\mathbb{R}}^{d}}f({\bf{x}})+\frac{2\eta}{\sqrt{m}}{\mathbb{E}}\left[\left\|{\bf V}^{0}-{\bf{1}}{\bar{\bf{v}}}^{0}\right\|\right]\leq 2\Delta_{f}$
and the requirement of above $\hat{\rho}$ can be satasfied by taking
$\hat{K}\geq\left\lceil\frac{1}{\sqrt{\gamma}}\log\frac{2\sqrt{2}\eta\sqrt{{m\sigma^{2}}/{b}+\sum_{i=1}^{m}\left\|\nabla f_{i}({\bf{x}}_{i}^{0})\right\|^{2}}}{\sqrt{m}\Delta_{f}}\right\rceil=\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{\gamma}}\right).$
Therefore, the term $\Delta_{\Phi}$ in the setting of $T$ presented in Theorem 1 can be replaced by $2\Delta_{f}$.
We then consider how to making the gradient small on each node. Specifically, we have
${\mathbb{E}}\left[\left\|\nabla f(\hat{\bf{x}}_{i})\right\|\right]=$ ${\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bf{x}}_{i}^{t})\right\|\right]$
$\leq$ ${\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bf{x}}_{i}^{t})-\nabla f({\bar{\bf{x}}}^{t})\right\|+\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right]$
$\leq$ ${\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\big{(}L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\big{)}\left\|{\bf{x}}_{i}^{t}-{\bar{\bf{x}}}^{t}\right\|+\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right]$
$\leq$ ${\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\big{(}L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\big{)}\left\|{\bf X}^{t}-{\bf{1}}{\bar{\bf{x}}}^{t}\right\|\right]+\frac{\epsilon}{2}$
$\leq$ ${\mathbb{E}}\left[\frac{1}{T}\sum_{t=0}^{T-1}\frac{\rho m\eta\big{(}L_{f}+L_{1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\big{)}}{1-\rho}\right]+\frac{\epsilon}{2}$
$=$ $\frac{\rho m\eta\left(L_{f}+{\mathbb{E}}\left[\frac{L_{1}}{T}\sum_{t=0}^{T-1}\left\|\nabla f({\bar{\bf{x}}}^{t})\right\|\right]\right)}{1-\rho}+\frac{\epsilon}{2}$
$\leq$ $\frac{\rho m\eta\big{(}L_{f}+L_{1}\epsilon/2\big{)}}{1-\rho}+\frac{\epsilon}{2}\leq\epsilon,$
where the first step is based on the triangle inequality; the second step is based on Proposition 1 and equation (18); the third step is based on Theorem 1; the fourth step is based on Lemma 2; the last line requires
$\rho\leq\frac{\epsilon}{2m\eta L_{f}+(m\eta L_{1}+1)\epsilon},$
which can be satisfies by taking $K=\tilde{\mathcal{O}}(\sqrt{1/\gamma})$.
In summary, the communication complexity is
$\hat{K}+TK=\tilde{\mathcal{O}}\left(\left(\frac{L_{f}}{\epsilon^{2}}+\frac{L_{1}}{\epsilon}\right)\frac{\Delta_{f}}{\sqrt{\gamma}}\right).$
Following equations (24), (25), and (26), the sample complexity on each node is
$Tb={\mathcal{O}}\left(\frac{1}{m}\left(\frac{L_{f}\sigma^{2}\Delta}{\epsilon^{4}}+\frac{\sigma^{2}}{\epsilon^{2}}+\frac{L_{1}^{3}\sigma^{2}\Delta}{L_{f}^{2}\epsilon}+\frac{L_{1}^{2}\sigma^{2}}{L_{f}^{2}}\right)\right),$
where the terms $\sigma^{2}/\epsilon^{2}$ and $L_{1}^{2}\sigma^{2}/L_{f}^{2}$ come from the initialization on ${\bf G}^{0}$ in line 3 of Algorithm 2. ∎
Appendix H The Proof of Proposition 3
Proof.
We can verify that the (second-order) derivatives of functions $f_{1}$, $f_{2}$, and $f$ have the form of
$f^{\prime}_{1}(x)=L\exp(Lx),\quad f^{\prime}_{2}(x)=-L\exp(-Lx),\quad\text{and}\quad f^{\prime}(x)=\frac{L\exp(Lx)-L\exp(-Lx)}{2}.$
Note that the function
$h(z)=\frac{|1-\exp(Lz)|}{L|z|}$
is increasing. Taking z=yxz=y-x, we have
$\frac{|1-\exp(L(y-x))|}{L|y-x|}=h(y-x)\leq h\left(\frac{1}{L_{1}}\right)=\frac{|1-\exp(L/L_{1})|}{L/L_{1}}$ (27)
for all $x,y\in{\mathbb{R}}$ with $|y-x|\in[-1/L_{1},1/L_{1}]$.
For the function $f_{1}$, we have
$|f^{\prime}_{1}(x)-f^{\prime}_{1}(y)|$
$=$ $|L\exp(Lx)-L\exp(Ly)|$
$=$ $L\exp(Lx)|1-\exp(L(y-x))|$
$\leq$ $L\exp(Lx)\cdot L_{1}|1-\exp(L/L_{1})||y-x|$
$=$ $L\exp(Lx)|y-x|L_{1}$
$=$ $L_{1}f_{1}^{\prime}(x)|x-y|,$
where the inequality is based on equation (27) and the last two steps are based on $f^{\prime}_{1}(x)=L\exp(Lx)$ and the setting $L_{1}=L/\log 2$. Therefore, we have shown that $f_{1}$ is $(0,L_{1})$-smooth. Similarly, we can also show that $f_{2}$ is $(0,L_{1})$-smooth.
Let $x=0$ and $y=(\log 2)/L$, which holds $|x-y|=1/L_{1}$. We also have
$|f^{\prime}(x)-f^{\prime}(y)|$
$=$ $\frac{L}{2}\left|\exp(Lx)-\exp(-Lx)-(\exp(Ly)-\exp(-Ly))\right|$
$=$ $\frac{L}{2}\left|\exp(0)-\exp(0)-\left(2-\frac{1}{2}\right)\right|=\frac{3L}{4}$
and
$L_{1}f^{\prime}(x)|y-x|=\frac{L_{1}L(\exp(Lx)-\exp(-Lx))|y-x|}{2}=0.$
Therefore, the inequality $|f^{\prime}(x)-f^{\prime}(y)|\leq L_{1}f^{\prime}(x)|y-x|$ does not hold, which means the function $f$ is not $(0,L_{1})$-smooth. ∎